GATE 2018 CE (CE2) – Question 29
A box contains 5 red balls and 1 white ball. The probability of consecutively drawing 3 red balls without replacement, to one decimal place, is ______.
Practise this question in The GATE Grind →
Show answer and explanation
Correct answer: 0.5 to 0.5
Explanation
Without replacement, multiply the conditional probabilities: $(5/6)(4/5)(3/4)=1/2=0.5$.